# Systematic Study on the $\alpha$-particle preformation factor in the theory of $\alpha$-decay based on the Tabular Prior-data Fitted Network (TabPFN)

### v1

@@ II.4 Preformation factor and nuclear structure information
The odd-even staggering (OES) effect is a well-established phenomenon manifested in various nuclear properties, including masses, nucleon separation energies, decay energies [yang2022new], and half-lives [yang2022examining]. It is therefore of interest to investigate whether this effect also manifests in $\alpha$-particle preformation factors. Figure 3 displays the preformation factors $P_{\alpha}^{\mathrm{TabPFN12}}$ computed using the TabPFN12 method for isotopic chains of Th, Pa, U, Np, and Pu. The results successfully reproduce the OES pattern. Specifically, for the even-$Z$ isotopic chains of Th, U, and Pu, $P_{\alpha}^{\mathrm{TabPFN12}}$ for odd-$N$ nuclei is systematically lower than for adjacent even-$N$ nuclei. This indicates that an unpaired neutron suppresses $\alpha$-cluster formation on the nuclear surface, thereby reducing $P_{\alpha}^{\mathrm{TabPFN12}}$. Similarly, comparing the odd-$Z$ Pa chain with the even-$Z$ Th chain reveals that, for the same neutron number $N$, $P_{\alpha}^{\mathrm{TabPFN12}}$ values for Pa isotopes are generally smaller, implying that an unpaired proton also hinders $\alpha$-particle formation. More systematically, the Th chain ($Z=90$) exhibits larger $P_{\alpha}^{\mathrm{TabPFN12}}$ values than the adjacent Pa chain ($Z=91$), and the Pu chain ($Z=94$) shows higher values than the neighboring Np chain ($Z=93$). At $N=144$, the preformation factor of U isotopes exceeds those of both the Th and Pu chains. These systematic trends, evident in Fig. 3, indicate that the $\alpha$-preformation factor increases with proton number after crossing the magic number $Z=82$, reaches a maximum around $Z=92$, and gradually decreases thereafter. The presence of unpaired nucleons reduces the probability of $\alpha$-cluster formation on the nuclear surface. Consequently, the behavior and evolution of $\alpha$-preformation factors provide valuable insights into nuclear structure.
Previous studies have commonly derived $\alpha$-particle preformation factors from empirical formulas. For comparison with the present TabPFN12-based results, the empirical formulations from Refs. [deng2021correlation] and [he2021improved] are adopted. The corresponding root-mean-square deviations $\sigma_{\mathrm{rms}}$ are summarized in Table II. A comparison demonstrates that the machine learning approach delivers performance comparable to or superior to that of empirical formulas. Specifically, relative to Ref. [he2021improved], the $\sigma_{\mathrm{rms}}$ values for the four nuclear categories decrease from 0.544, 0.632, 0.564, and 0.610 to 0.153, 0.254, 0.220, and 0.249, respectively. When compared to Ref. [deng2021correlation], the corresponding $\sigma_{\mathrm{rms}}$ values are reduced from 0.548, 0.713, 0.655, and 0.646 to the same set of improved values. The overall $\sigma_{\mathrm{rms}}$ is reduced from 0.584 and 0.637 to 0.211, corresponding to improvements of 63.8% and 66.8%, respectively.
@@
nuclei type TabPFN12 ref. [he2021improved] ref. [deng2021correlation]
@@ II.5 Extrapolation capability of TabPFN12 neural network
The $\alpha$-decay half-life is an experimentally measurable quantity. Theoretical half-lives incorporating the $\alpha$-particle preformation probability are expected to show improved agreement with experimental data. To assess the performance of the TabPFN12-derived preformation factors, we compute them for 41 newly evaluated nuclei with $Z>104$ in NUBASE2020—an update from the 2016 evaluation—and compare them with predictions from the empirical formula of Ref. [deng2021correlation]. The corresponding $\alpha$-decay energies are taken from the AME2020 atomic mass evaluation. Detailed results are summarized in Table III. The $P_{\alpha}$ values from the TabPFN12 approach are largely consistent with the empirical systematics, indicating that the model can reliably extrapolate preformation factors to unexplored regions of the nuclear chart.
Furthermore, $\alpha$-decay half-lives are computed by incorporating the TabPFN12 preformation factor and compared with experimental values. Fig. 4 shows the deviations $\log_{10}(T^{\mathrm{exp}}{1/2}/T^{\mathrm{cal}}{1/2})$ between experimental half-lives and CPPM calculations for the 498 nuclei and the additional 41 nuclei, considering cases both with and without the preformation factor. Inclusion of the preformation factor reduces the root-mean-square (rms) deviation for the 498 nuclei from 2.028 to 0.211—an improvement of 89.5%. For the 41 test nuclei, the rms deviation drops from 2.920 to 0.771, corresponding to a 73.5% improvement. The significantly improved accuracy of the half-life predictions validates the reliability of the preformation factors obtained from the TabPFN12 method.
Table 3: A comparison of $\alpha$-particle preformation factors between TabPFN predictions and empirical formulations from Refs. [deng2021correlation, he2021improved] is presented for 41 newly evaluated nuclei in the NUBASE2020 database [kondev2021nubase2020]. The quantities $P^{\text{Deng}}_{\alpha}$ and $P^{\text{He}}_{\alpha}$ correspond to preformation factors calculated using the empirical relations in Refs. [deng2021correlation] and [he2021improved], respectively, while $P^{\text{TabPFN12}}_{\alpha}$ denotes the values predicted by the TabPFN12 model.
@@
Figure 4: Upper panel: Differences between experimental and theoretical half-lives for 498 nuclei within the CPPM framework, showing results both with and without inclusion of the $P_{\alpha}^{\mathrm{TabPFN12}}$ preformation factor. Lower panel: Corresponding deviations for the additional set of 41 nuclei.
Table 4: Predicted $\alpha$-decay half-lives for superheavy nuclei with Z=117120Z=117-120, where decay energies (in MeV) are adopted from Ref. [ma2019basic]. Half-lives are presented in units of seconds.
@@
Accurate prediction of $\alpha$-decay half-lives is crucial for both the synthesis of superheavy elements and the understanding of their structural properties. The thoroughly trained TabPFN12 model demonstrates reliable predictive capability for $\alpha$-particle preformation factors. This model is therefore employed to predict the preformation factors for superheavy nuclei with Z=117Z=117–120, utilizing $\alpha$-decay energies $Q_{\alpha}$ from Ref. [ma2019basic]. Figure 5(a) presents the predicted preformation factors for $Z=117$ and 118 isotopes, revealing a distinct odd-even staggering (OES) effect. For instance, at $N=182$, $P_{\alpha}^{\mathrm{TabPFN12}}$ for $Z=118$ is larger than for its $Z=117$ isotone, decreasing from $\sim$0.0088 to 0.0044—a reduction of 50%. Similarly, within the $Z=118$ chain, $P_{\alpha}^{\mathrm{TabPFN12}}$ for $N=182$ exceeds that for $N=183$ (0.0039). A similar systematic behavior is observed for $Z=119$ and 120 nuclei, as shown in Fig. 5(b). These results indicate that both unpaired protons and unpaired neutrons inhibit $\alpha$-particle preformation. Furthermore, the $P_{\alpha}^{\mathrm{TabPFN12}}$ values for the even-$Z$ chains (120, 118) are systematically larger than for the odd-$Z$ chains (119, 117), highlighting the key role of proton pairing correlations in $\alpha$-cluster preformation.
@@
The predicted preformation factors exhibit clear dependencies on nuclear structure, including pronounced odd-even staggering and shell closures at $Z=82$ and $N=126$. The TabPFN12 model successfully captures the linear correlation between $\log_{10}P_{\alpha}$ and $Q_{\alpha}^{-1/2}$, effectively extending the Geiger-Nuttall systematics to preformation factors. The results confirm that unpaired nucleons significantly inhibit $\alpha$-cluster formation, as evidenced by systematically lower preformation factors in odd-$A$ nuclei compared to their even-even neighbors.
Incorporated into the CPPM framework, the TabPFN12-predicted preformation factors substantially improve $\alpha$-decay half-life calculations, reducing the RMS deviation from 2.028 to 0.211 (89.5% improvement) for the training set and from 2.920 to 0.771 (73.5% improvement) for 41 newly evaluated superheavy nuclei. The model’s extrapolation capability is further validated by predictions for Z=117Z=117–120 isotopic chains, which reveal a sharp decrease in half-lives from $N=184$ to $N=186$. This rapid change suggests that $N=184$ may constitute a new neutron magic number in the superheavy region.
@@ IV Acknowledgements
This work is supported by Yunnan Provincial Science Foundation Project (No. 202501AT070067), Yunnan Provincial Xing Dian Talent Support Program (Young Talents Special Program, No. XDYC-QNRC-2023-0162), Kunming University Talent Introduction Research Project (No. YJL24019), Yunnan Provincial Department of Education Scientific Research Fund Project (No. 2025Y1042 and 2025Y1055), the Program for Frontier Research Team of Kunming University 2023, and National Natural Science Foundation of China (No. 12063006), the Special Basic Cooperative Research Programs of Yunnan Provincial Undergraduate Universities’ Association (grant NO. 202101BA070001-144).

### v2

@@ II.4 Preformation factor and nuclear structure information
The odd-even staggering (OES) effect is a well-established phenomenon manifested in various nuclear properties, including masses, nucleon separation energies, decay energies [62], and half-lives [61]. It is therefore of interest to investigate whether this effect also manifests in $\alpha$-particle preformation factors. Figure 4 displays the preformation factors $P_{\alpha}^{\mathrm{TabPFN12}}$ computed using the TabPFN12 method for isotopic chains of Th, Pa, U, Np, and Pu. The results successfully reproduce the OES pattern. Specifically, for the even-$Z$ isotopic chains of Th, U, and Pu, $P_{\alpha}^{\mathrm{TabPFN12}}$ for odd-$N$ nuclei is systematically lower than for adjacent even-$N$ nuclei. This indicates that an unpaired neutron suppresses $\alpha$-cluster formation on the nuclear surface, thereby reducing $P_{\alpha}^{\mathrm{TabPFN12}}$. Similarly, comparing the odd-$Z$ Pa chain with the even-$Z$ Th chain reveals that, for the same neutron number $N$, $P_{\alpha}^{\mathrm{TabPFN12}}$ values for Pa isotopes are generally smaller, implying that an unpaired proton also hinders $\alpha$-particle formation. More systematically, the Th chain ($Z=90$) exhibits larger $P_{\alpha}^{\mathrm{TabPFN12}}$ values than the adjacent Pa chain ($Z=91$), and the Pu chain ($Z=94$) shows higher values than the neighboring Np chain ($Z=93$). At $N=144$, the preformation factor of U isotopes exceeds those of both the Th and Pu chains. These systematic trends, evident in Fig. 4, indicate that the $\alpha$-preformation factor increases with proton number after crossing the magic number $Z=82$, reaches a maximum around $Z=92$, and gradually decreases thereafter. The presence of unpaired nucleons reduces the probability of $\alpha$-cluster formation on the nuclear surface. Consequently, the behavior and evolution of $\alpha$-preformation factors provide valuable insights into nuclear structure.
Previous studies have commonly derived $\alpha$-particle preformation factors from empirical formulas. For comparison with the present TabPFN12-based results, the empirical formulations from Refs. [15] and [24] are adopted. The corresponding root-mean-square deviations $\sigma_{\mathrm{rms}}$ are summarized in Table II. A comparison demonstrates that the machine learning approach delivers performance comparable to or superior to that of empirical formulas. Specifically, relative to Ref. [24], the $\sigma_{\mathrm{rms}}$ values for the four nuclear categories decrease from 0.544, 0.632, 0.564, and 0.610 to 0.153, 0.254, 0.220, and 0.249, respectively. When compared to Ref. [15], the corresponding $\sigma_{\mathrm{rms}}$ values are reduced from 0.548, 0.713, 0.655, and 0.646 to the same set of improved values. The overall $\sigma_{\mathrm{rms}}$ is reduced from 0.584 and 0.637 to 0.211, corresponding to improvements of 63.8% and 66.8%, respectively.
@@
nuclei type TabPFN12 ref. [24] ref. [15]
@@ II.5 Extrapolation capability of TabPFN12 neural network
The $\alpha$-decay half-life is an experimentally measurable quantity. Theoretical half-lives incorporating the $\alpha$-particle preformation probability are expected to show improved agreement with experimental data. To assess the performance of the TabPFN12-derived preformation factors, we compute them for 41 newly evaluated nuclei with $Z>104$ in NUBASE2020—an update from the 2016 evaluation—and compare them with predictions from the empirical formula of Ref. [15]. The corresponding $\alpha$-decay energies are taken from the AME2020 atomic mass evaluation. Detailed results are summarized in Table III. The $P_{\alpha}$ values from the TabPFN12 approach are largely consistent with the empirical systematics, indicating that the model can reliably extrapolate preformation factors to unexplored regions of the nuclear chart.
Furthermore, $\alpha$-decay half-lives are calculated by incorporating the TabPFN12 preformation factors and compared with experimental data. Figure 5 shows the deviations $\log_{10}(T^{\mathrm{exp}}_{1/2})-\log_{10}(T^{\mathrm{cal}}_{1/2})$ between experimental and calculated half-lives. Figure 5 (a) presents the results for 498 nuclei within the CPPM framework. Without including the preformation factor, the calculated half-lives show large systematic deviations from the experimental values, yielding a root-mean-square deviation $\sigma_{\mathrm{rms}}=2.028$. After incorporating the TabPFN12 preformation factor, the deviations become concentrated around zero and the rms deviation is significantly reduced to $\sigma_{\mathrm{rms}}=0.211$, corresponding to an improvement of about 89.5%. To further test the predictive capability of the model, the same procedure is applied to an additional set of 41 nuclei, as shown in Fig. 5 (b). A similar trend is observed: the rms deviation decreases from 2.920 to 0.771 when the TabPFN12 preformation factor is included, indicating that the model retains good predictive performance for nuclei outside the training dataset. For comparison, Fig. 5 (c) shows the deviations obtained using a deformed Coulomb plus Woods–Saxon potential for the same 498 nuclei. The inclusion of the TabPFN12 preformation factor again significantly reduces the deviations, with the rms value decreasing from 0.954 to 0.232. Although this value is slightly larger than that obtained within the CPPM framework ($\sigma_{\mathrm{rms}}=0.211$), the calculated half-lives remain well clustered around the experimental values. This indicates that the TabPFN12-derived preformation factors provide a robust improvement to $\alpha$-decay half-life predictions for both spherical and deformed potential descriptions.
Table 3: A comparison of $\alpha$-particle preformation factors between TabPFN predictions and empirical formulations from Refs. [15, 24] is presented for 41 newly evaluated nuclei in the NUBASE2020 database [33]. The quantities $P^{\text{Deng}}_{\alpha}$ and $P^{\text{He}}_{\alpha}$ correspond to preformation factors calculated using the empirical relations in Refs. [15] and [24], respectively, while $P^{\text{TabPFN12}}_{\alpha}$ denotes the values predicted by the TabPFN12 model.
@@
Figure 5: Deviations between experimental and calculated $\alpha$-decay half-lives. (a) Results for 498 nuclei obtained within the CPPM framework. (b) Predictions for an additional set of 41 nuclei. (c) Results for 498 nuclei calculated using the deformed Coulomb plus Woods–Saxon potential. Blue stars denote calculations including the TabPFN12 preformation factor, while red circles correspond to results without the preformation factor.
Table 4: Predicted $\alpha$-decay half-lives for superheavy nuclei with Z=117120Z=117-120, where decay energies (in MeV) are adopted from Ref. [36]. Half-lives are presented in units of seconds.
@@
Accurate prediction of $\alpha$-decay half-lives is crucial for both the synthesis of superheavy elements and the understanding of their structural properties. The thoroughly trained TabPFN12 model demonstrates reliable predictive capability for $\alpha$-particle preformation factors. This model is therefore employed to predict the preformation factors for superheavy nuclei with Z=117Z=117–120, utilizing $\alpha$-decay energies $Q_{\alpha}$ from Ref. [36]. Figure 6 (a) presents the predicted preformation factors for $Z=117$ and 118 isotopes, revealing a distinct odd-even staggering (OES) effect. For instance, at $N=182$, $P_{\alpha}^{\mathrm{TabPFN12}}$ for $Z=118$ is larger than for its $Z=117$ isotone, decreasing from $\sim$0.0088 to 0.0044—a reduction of 50%. Similarly, within the $Z=118$ chain, $P_{\alpha}^{\mathrm{TabPFN12}}$ for $N=182$ exceeds that for $N=183$ (0.0039). A similar systematic behavior is observed for $Z=119$ and 120 nuclei, as shown in Fig. 6 (b). These results indicate that both unpaired protons and unpaired neutrons inhibit $\alpha$-particle preformation. Furthermore, the $P_{\alpha}^{\mathrm{TabPFN12}}$ values for the even-$Z$ chains (120, 118) are systematically larger than for the odd-$Z$ chains (119, 117), highlighting the key role of proton pairing correlations in $\alpha$-cluster preformation.
@@
The predicted preformation factors exhibit clear dependencies on nuclear structure, including pronounced odd-even staggering and shell closures at $Z=82$ and $N=126$. The TabPFN12 model successfully captures the linear correlation between $\log_{10}P_{\alpha}$ and $Q_{\alpha}^{-1/2}$, effectively extending the Geiger-Nuttall systematics to preformation factors. A clear negative correlation between $\log_{10}P_{\alpha}$ and the fragmentation potential $V_{\mathrm{frag}}$ is also observed. The results confirm that unpaired nucleons significantly inhibit $\alpha$-cluster formation, as evidenced by systematically lower preformation factors in odd-$A$ nuclei compared to their even-even neighbors.
Incorporated into the CPPM framework, the TabPFN12-predicted preformation factors substantially improve $\alpha$-decay half-life calculations, reducing the RMS deviation from 2.028 to 0.211 (89.5% improvement) for the training set and from 2.920 to 0.771 (73.5% improvement) for 41 newly evaluated superheavy nuclei. A similar improvement is also obtained when a deformed Coulomb plus Woods–Saxon potential is adopted, yielding an rms deviation of $\sigma_{\mathrm{rms}}=0.232$, which remains close to the value obtained within the CPPM framework and demonstrates the robustness of the TabPFN-derived preformation factors against different potential descriptions. The model’s extrapolation capability is further validated by predictions for Z=117Z=117–120 isotopic chains, which reveal a sharp decrease in half-lives from $N=184$ to $N=186$. This rapid change suggests that $N=184$ may constitute a new neutron magic number in the superheavy region.
@@ IV Acknowledgements
This work is supported by Yunnan Provincial Science Foundation Project (No. 202501AT070067), Yunnan Provincial Xing Dian Talent Support Program (Young Talents Special Program, No. XDYC-QNRC-2023-0162), Kunming University Talent Introduction Research Project (No. YJL24019), Yunnan Provincial Department of Education Scientific Research Fund Project (No. 2025Y1042 and 2025Y1055), the Program for Frontier Research Team of Kunming University 2023, National Natural Science Foundation of China (No. 12063006), the Special Basic Cooperative Research Programs of Yunnan Provincial Undergraduate Universities’ Association (grant NO. 202101BA070001-144), and Xing Dingyu Academician Workstation of Yunnan Province (No. 202605AF350035).